# Language models’ activations linearly encode training-order recency

### v1

@@ Abstract
We show that language models’ activations linearly encode when information was learned during training. Our setup involves creating a model with a known training order by sequentially fine-tuning Llama-3.2-1B on six disjoint but otherwise similar datasets about named entities. We find that the average activations of test samples for the six training datasets encode the training order: when projected into a 2D subspace, these centroids are arranged exactly in the order of training and lie on a straight line. Further, we show that linear probes can accurately ($\sim$90%) distinguish “early” vs. “late” entities, generalizing to entities unseen during the probes’ own training. The model can also be fine-tuned to explicitly report an unseen entity’s training stage ($\sim$80% accuracy). Interestingly, this temporal signal does not seem attributable to simple differences in activation magnitudes, losses, or model confidence. Our paper demonstrates that models are capable of differentiating information by its acquisition time, and carries significant implications for how they might manage conflicting data and respond to knowledge modifications.
@@ 1 Introduction
We show that a single linear direction in activation space captures training order across these six fine-tuning stages. Specifically, we take test samples’ activations and average them according to the stage at which the corresponding entity was encountered in training – thus giving us six test activation centroids. When orthonormally projected into a 2D subspace, these six centroids lie on a straight line along a “training-order recency” direction which is consistent across independent fine-tuning runs with different datasets (Figure 1).
Could models access this information, or are activation distributions too broad for models to reliably know when they learned about a given entity? Linear probes reach >>90% accuracy at distinguishing entities introduced early in fine-tuning (those in $D_{1}$) from those introduced later ($D_{6}$). To ensure probes learn a general pattern rather than memorising early vs. late entities, we train them on one subset of entities that were used to fine-tune the LLM, and evaluate on a disjoint subset. Similarly, we can fine-tune the model to answer questions like “Which training stage is this alias from?”. The resulting model reaches $\sim$80% accuracy on aliases it never saw in this auxiliary fine-tune, confirming that fine-tuning makes this information directly accessible.
Remarkably, the training-order encoding is rather persistent: after 30 additional fine-tuning epochs on joint and shuffled data from the initial six stages—where no training-order distinction is reinforced—probe accuracy only decays to $\sim$63%, well above the 50% chance level. Additionally, our careful distribution balancing experiments show that this temporal encoding is not a trivial artifact of training or validation loss, activation norms, or the model’s confidence.
@@
(1) We provide the first empirical evidence that training‑order information is linearly encoded in LLM activations; (2) we show that this encoding generalizes across independent fine-tuning runs with different kinds of training data; (3) we demonstrate this encoding’s persistence under further joint training and rule out several simple explanations; and (4) we verify that the model can exploit this information when fine-tuned to do so.
[…]
@@ 3.1 Training order is linearly encoded in a subspace
While the activations’ centroids line up in Figure 1, are their full distributions distinct enough for it to be feasible to tell a given entity’s training stage? Linear probes trained to distinguish entities from two specific stages perform best when spanning broad temporal intervals: notably, a probe distinguishing the earliest from the latest stages, $D_{1}$ vs. $D_{6}$, achieves over 90% accuracy (see Figure 2). This high accuracy occurs despite the probe-training and evaluation sets containing entirely disjoint entities, showing that the probes generalize beyond specific memorized examples. Probes trained on narrower intervals such as adjacent stages perform worse: robust classification is best achieved using broader temporal comparisons.
@@ 3.2 This direction seems to track exposure recency
Figure 3a shows intermediate model checkpoints after each of the six training stages. For each checkpoint, activation centroids line up as “never seen” $\to$ “seen in the past” $\to$ “seen recently” when projected onto the same x-axis as in Figure 1.
Figure 3: Most recently seen data is on one side, never-seen data is on the other. We track centroids’ evolution over the course of training by plotting them for checkpoints after each training stage. Distinct marker shapes correspond to different checkpoints; hollow markers denote centroids whose corresponding training dataset was not yet trained on. After each training stage, that stage’s centroid ends up furthest to the right. Centroids for the not-yet-seen datasets are on the very left, giving a “never seen” to “seen most recently” axis. a) The projections’ x-axis is identical to that in Figure 1. b) Same centroids as in a) but projected onto a “seen” vs. “never-seen” x-axis. This axis also orders the centroids according to the training order.
Interestingly, for checkpoints after stage 4 and later, the earliest stage’s training dataset $D_{1}$ is indistinguishable from not-yet-seen datasets like $D_{5}$ or $D_{6}$. We found such collapse to be an artifact of this specific projection: Figure 3b shows a different projection where never-seen datasets’ centroids are always clearly distinguishable from centroids for previously seen data. The x-axis there is computed as the average difference between not-yet-seen activations’ centroids ($D_{6}$ for all checkpoints before the 6th stage), and the centroids corresponding to $D_{1}$ (the dataset that all these checkpoints were already trained on). This “never seen in training” vs. “seen” axis also orders the centroids according to the ground truth training order, implying that the previously observed ability of probes to distinguish entities that were and were not present in training (Ferrando et al., 2024) might be due to the richer temporal encoding – or further still, both phenomena might be reflections of how models encode some notion of confidence.
Figure 4: a) Re-exposing one of the six datasets as an additional 7th finetuning stage moves the re-exposed dataset’s centroid to the very right, reflecting this dataset as the most recent one. Interestingly, re-exposure fine-tuning sometimes reverses the order of the intermediate centroids, e.g. when re-exposing stage 2. b) Additional training moves centroids further right: training one stage for 15 epochs instead of the usual 5 widens the gap from the previous stage’s centroid. The projections’ x-axis in both subplots is identical to that in Figure 1; the orthogonal y-axis is unique to each subplot.
@@ 3.5 Models can explicitly report training stages
The fine-tuned model achieves 79.8% accuracy on held-out “probe-test” aliases, far exceeding the 50% random baseline and approaching the $\sim 90\%$ accuracy of linear probes. This demonstrates that the training-order information encoded in activations is not merely detectable by external analysis but is actively accessible to the model’s own computations. While we cannot claim that models use this signal during standard inference, establishing that they can access temporal information when needed is the key finding. The capability exists, and if distinguishing training stages helps achieve lower loss—perhaps through strategic behavior or “playing the training game”—models might spontaneously learn to leverage this latent information.
@@ 6 Discussion
Our experiments are restricted to fine-tuning relatively small models ($\leq$8B parameters) on two variants of an essentially toy dataset. Several aspects of our findings’ generality remain unexplored. Our work focuses exclusively on language models, and it’s an open question whether analogous temporal encoding mechanisms exist in other architectures and modalities such as vision. And in the language setting, would the training order encoding emerge when training from scratch or using more naturalistic data? Finally, while our experiments in §4 show that several simple explanations are insufficient to explain the training-order encoding, it is still unclear what exactly underpins the phenomenon.
@@
If the training mixture is “bursty” across documents – e.g. documents on the same topic are often close to each other in the order of training, even if they are not in the same context – it’d help to remember which examples are more recent. Specifically, given such structure of the training data a model might achieve lower loss by retrieving information related to recently-seen training data over data seen further in the past – thus the loss might encourage keeping track of the training order. This is analogous to a burstiness-based explanation for in-context learning given by Chan et al. (2022). However, for this mechanism to explain our results, the training order would have needed to be non-random during pre-training, making the explanation less plausible. We believe this condition is needed since the training-order direction is likely already present in the pre-trained model, as indicated by the different fine-tuning runs’ lines being parallel in Figure 1.
Future work could explore whether it is possible to learn about the order of pre-training data using the training-order axis discovered during fine-tuning. It would also be interesting to study how conflicting training data, e.g. contradicting statements present in different training stages, is reflected on the training-order axis. Another promising direction is studying whether models can in fact use training-order information to resist modification, and how. If several beliefs are all inserted at the same time during fine-tuning as done by Wang et al. (2025), could the model identify multiple such implants upon finding one at test-time? If such identification is possible, we could develop techniques for making the inserted belief look like it was learned a long time ago. Relatedly, could models use training-order information to resist modification at training time, like in the Alignment Faking scenarios from Greenblatt et al. (2024)? Specifically, could models learn about what they’re being trained to do by accessing information they were recently trained on – and could the training-order axis help with that?
@@ Conclusion.
Language models linearly encode when they learned what they know—a persistent timestamp that generalizes across independent fine-tuning runs and withstands prolonged training on shuffled data. Models can access this information when fine-tuned to do so, achieving 80% accuracy at reporting their own training history. This discovery reveals a fundamental property of how neural networks organize knowledge and opens up a number of promising research directions.
@@ Acknowledgements
We thank Ryan Greenblatt, Ekdeep Singh Lubana, Stewart Slocum, Stefan Heimersheim, Bruno Mlodozeniec, and Fabien Roger for feedback and helpful discussions.
@@ Appendix C Additional results
Figure 11: Projections of activations onto the direction of probes distinguishing $D_{1}$ vs. $D_{6}$ for the model after six-stage sequential fine-tuning. Despite never seeing $D_{2},...,D_{5}$ during training, the probes correctly order all six stages by recency (dotted lines = means). Layer 13/16, last token position (token “:” in Figure 2b). Left: logistic regression probe, right: diffmean probe. Visually, the direction of the logistic regression probe better separates the different stages’ activations. Note that the axes are scaled differently – specifically, the diffmean plot’s axis is scaled s.t. $D_{1}$ centroid is at -1 and $D_{6}$ centroid has value 1.
@@ E.2 Results
For certain token positions, balanced probes perform noticeably worse than random controls (see Figures 1516 and 17). However, while these simple distribution differences contribute to probe accuracy in specific circumstances, they cannot fully account for the phenomenon, since the recency encoding persists even when controlling for these differences.

### v2

# Fresh in memory: Training-order recency is linearly encoded in language model activations

@@ Abstract
We show that language models’ activations linearly encode when information was learned during training. Our setup involves creating a model with a known training order by sequentially fine-tuning Llama-3.2-1B on six disjoint but otherwise similar datasets about named entities. We find that the average activations of test samples corresponding to the six training datasets encode the training order: when projected into a 2D subspace, these centroids are arranged exactly in the order of training and lie on a straight line. Further, we show that linear probes can accurately ($\sim$90%) distinguish “early” vs. “late” entities, generalizing to entities unseen during the probes’ own training. The model can also be fine-tuned to explicitly report an unseen entity’s training stage ($\sim$80% accuracy). Interestingly, the training-order encoding does not seem attributable to simple differences in activation magnitudes, losses, or model confidence. Our paper demonstrates that models are capable of differentiating information by its acquisition time, and carries significant implications for how they might manage conflicting data and respond to knowledge modifications.
@@ 1 Introduction
We show that a single linear direction in activation space captures training order across these six fine-tuning stages. Specifically, we take test samples’ activations and average them according to the stage at which the corresponding entity was encountered in training – thus giving us six test activation centroids. When projected into a 2D subspace, these six centroids end up arranged exactly in the order of training and lie on a straight line – a “training-order recency” direction consistent across independent fine-tuning runs with different datasets (Figure 1).
Could models access this information, or are activation distributions too broad for models to know when they learned about a given entity? Linear probes reach >>90% accuracy at distinguishing entities introduced early in fine-tuning (those in $D_{1}$) from those introduced later ($D_{6}$).
To ensure probes learn a general pattern rather than memorising early vs. late entities, we train them on one subset of entities that were used to fine-tune the LLM, and evaluate on a disjoint subset. Similarly, we can fine-tune the model to answer questions like “Which training stage is this alias from?”. The resulting model reaches $\sim$80% accuracy on aliases it never saw in this auxiliary fine-tune, confirming that fine-tuning makes this information directly accessible.
Remarkably, the training-order encoding is rather persistent: after 30 additional fine-tuning epochs on joint and shuffled data from the initial six stages – where no training-order distinction is reinforced – probe accuracy only decays to $\sim$63%, well above the 50% chance level. We find that the training-order encoding connects to binary knowledge detection (Ferrando et al., 2024): the “seen-in-training vs. unseen” axis also recovers the full training sequence, suggesting a shared underlying mechanism. Additionally, our careful distribution balancing experiments show that this temporal encoding is not a trivial artifact of training or validation loss, activation norms, or the model’s confidence.
@@
  1. 1.
    We provide the first evidence that training-order information is linearly encoded in LLM activations. This encoding generalizes across independent fine-tuning runs with different kinds of data.
  2. 2.
    We show the encoding tracks recency (re-training on earlier stages’ data moves corresponding centroids to the “most recent” position) and persists through additional mixed-data training.
  3. 3.
    We rule out simple explanations: the encoding is not fully explained by activation statistics (e.g. magnitudes) or measures of the model’s confidence.
  4. 4.
    We verify that the model can exploit training-order information when fine-tuned to do so.
@@ 3.1 Training order is linearly encoded in a subspace
While the activations’ centroids line up in Figure 1, are their full distributions distinct enough for it to be feasible to tell a given entity’s training stage? Linear probes trained to distinguish entities from two specific stages perform best when spanning broad temporal intervals: notably, a probe distinguishing the earliest from the latest stages achieves over 90% accuracy (see Figure 2). This high accuracy occurs despite the probe-training and evaluation sets containing entirely disjoint entities, showing that the probes generalize beyond specific memorized examples. Probes trained on narrower intervals such as adjacent stages perform worse. Appendix Figures 910 and 11 visualize activation distributions, which appear to overlap substantially despite the high probe accuracy.
@@ 3.2 This direction seems to track exposure recency
Figure 3a shows intermediate model checkpoints after each training stage. For each checkpoint, activation centroids line up as “never seen” $\to$ “seen in the past” $\to$ “seen recently” when projected onto the same x-axis as in Figure 1. Interestingly, for checkpoints after stage 4 and later, the earliest stage’s training dataset $D_{1}$ is indistinguishable from not-yet-seen datasets like $D_{5}$ or $D_{6}$. We found such collapse to be an artifact of this specific projection: Figure 3b shows a different projection where never-seen datasets’ centroids are always clearly distinguishable from centroids for previously seen data.
Figure 3: Most recently seen data is on one side, never-seen data is on the other. We track centroids’ evolution during training by plotting them for checkpoints after each training stage. Marker shapes denote different checkpoints; hollow markers indicate centroids whose corresponding training dataset was not yet trained on. After each training stage, that stage’s centroid ends up furthest to the right. Centroids for the not-yet-trained-on datasets are on the very left, giving a “never seen” to “seen most recently” axis. a) The projections’ x-axis is identical to that in Figure 1. b) Same centroids projected onto a “seen” vs. “never-seen” x-axis; this also orders the centroids according to the training order.
The x-axis in Figure 3b is computed as the average difference between not-yet-seen activations’ centroids ($D_{6}$ for all checkpoints before the 6th stage), and the centroids corresponding to $D_{1}$ (the dataset that all these checkpoints were already trained on). This “never seen in training” vs. “seen” axis also orders the centroids according to the ground truth training order, implying that the previously observed ability of probes to distinguish entities that were and were not present in training (Ferrando et al., 2024) might be due to the richer temporal encoding – or further still, both phenomena might be reflections of how models encode some notion of confidence.
Figure 4: Encoding tracks training recency and reflects the amount of training. a) Re-exposing one of the six datasets as an additional 7th fine-tuning stage moves the re-exposed dataset’s centroid to the very right, reflecting this dataset as the most recent one. Interestingly, re-exposure fine-tuning sometimes reverses the order of the intermediate centroids, e.g. when re-exposing stage 2. b) Additional training moves centroids further right: training one stage for 15 epochs instead of the usual 5 widens the gap from the previous stage’s centroid. The projections’ x-axis in both subplots is identical to that in Figure 1; the orthogonal y-axis is unique to each subplot.
@@ 3.5 Models can explicitly report training stages
The fine-tuned model achieves 79.8% accuracy on held-out “probe-test” aliases, showing that the training-order information encoded in the activations is not only detectable by external analysis but is also actively accessible to the model’s own computations. While we cannot claim that models use this signal during standard inference, establishing that they can access training-order information when needed is an important observation on its own. Given this capability, if distinguishing training stages can help models achieve lower loss – perhaps through strategic behavior or “playing the training game” – models might spontaneously learn to leverage this latent information.
@@ 6 Discussion
Our experiments are restricted to fine-tuning relatively small models ($\leq$8B parameters) on two variants of a relatively toy dataset. Several aspects of our findings’ generality remain unexplored. Our work focuses exclusively on language models, and it’s an open question whether an analogous training-order encoding can be found in other architectures and modalities such as vision. And in the language setting, would this encoding emerge when training from scratch? Additionally, can LLMs ever access this information without explicit fine-tuning like in §3.5? Finally, while we show that several simple explanations are insufficient to explain the training-order encoding (§4), it is still unclear what exactly underpins the phenomenon.
@@
If the training mixture is “bursty” across documents – e.g. documents on the same topic are often close to each other in the order of training, even if they are not in the same context – it’d help to remember which examples are more recent. Specifically, given such structure of the training data a model might achieve lower loss by retrieving information related to recently-seen training data over data seen further in the past – thus the loss might encourage keeping track of the training order. This is analogous to a burstiness-based explanation for in-context learning given by Chan et al. (2022). However, for this mechanism to explain our results, the training order would have needed to be non-random during pre-training, making the explanation less plausible. We believe this condition is needed since the training-order direction is likely already present in the pre-trained model, as indicated by the different fine-tuning runs’ lines being roughly parallel in Figure 1.
Future work could explore whether it is possible to learn about the order of pre-training data using the training-order axis discovered during fine-tuning. It would also be interesting to study how conflicting training data, e.g. contradicting statements present in different training stages, is reflected on the training-order axis. Another promising direction is studying whether models can in fact use training-order information to resist modification, and how. If several beliefs are all inserted at the same time during fine-tuning as done by Wang et al. (2025), could the model identify multiple such implants upon finding one at test-time? If such identification is possible, we could develop techniques for making the inserted belief look like it was learned a long time ago. Relatedly, could models use training-order information to resist modification at training time, like in the Alignment Faking scenarios from Greenblatt et al. (2024)? Specifically, could models learn about what they’re being trained to do by accessing information they were recently trained on – e.g. realize that the last few training steps changed the model to favor a particular answer style – and could the training-order axis help with that?
@@ Conclusion.
Language models linearly encode when they learned what they know – a timestamp that generalizes across independent fine-tuning runs and withstands prolonged training on shuffled data. Models can access this information when fine-tuned to do so, achieving 80% accuracy at reporting their own training history. This discovery hints at a fundamental property of how neural networks organize knowledge, and opens up a number of promising research directions.
@@ Acknowledgements
We thank Ryan Greenblatt, Ekdeep Singh Lubana, Stewart Slocum, Stefan Heimersheim, Bruno Mlodozeniec, Neel Alex, and Fabien Roger for feedback and helpful discussions.
@@ Appendix C Additional results
Figure 10: Plot identical to Figure 1 but including activations’ 68% KDE for one of the runs.
Figure 11: Projections of activations onto the direction of probes distinguishing $D_{1}$ vs. $D_{6}$ for the model after six-stage sequential fine-tuning. Left: diffmean axis, right: normal vector to the decision boundary of the logistic regression probe. Visually, the logistic regression probe’s axis better separates activations of the the different stages’ test data. Note that the axes are scaled differently – specifically, here the diffmean plot’s axis is scaled s.t. $c_{1}$ is at -1 and $c_{6}$ has value 1. Layer 13/16, last token position.
@@ E.2 Results
For certain token positions, balanced probes perform noticeably worse than random controls (see Figures 1617 and 18). However, while these simple distribution differences contribute to probe accuracy in specific circumstances, they cannot fully account for the phenomenon, since the recency encoding persists even when controlling for these differences.